# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v3

@@ Abstract
Project page: ramyamut.github.io/e3-pose/
@@ 1 Introduction
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [billard2019], autonomous driving [chiu2021], and image-guided surgery [yaniv2006]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology [glenn2009, gholipour2014] (Appendix A). Due to inter-slice fetal motion, oblique (i.e., misaligned with the target anatomical orientation) slices and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack (Fig. 1) [malamateniou2013, xu2023, white2010]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The pose estimated from each navigator volume can be used to adaptively adjust the imaging plane of the next 2D slice in real time (Fig. 1) [hess2011, tisdall2012, gagoski2016]. Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and low SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) diagnostic 2D slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates a left-right symmetry, introducing pose ambiguities [zhan2013]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@ 4.2 Rotation estimation
Let $G_{\text{symm}}\!=\!\{g_{s}\in\text{O(3)}\!:\!f\!\circ\!g_{s}^{-1}\!=\!f\}$ be the object symmetry group [saxena2007]. To handle object symmetries, we aim to parametrize $h(R)$ such that $\rho_{h}(g_{s})h(R)\!=\!h(R)$ for all $g_{s}\!\in\!G_{\text{symm}}$ and $R\!\in\!\text{SO(3)}$; and $h$ remains continuous over the quotient group $H\!=\!\text{O(3)}/G_{\text{symm}}$. Here, we consider the case $S\!=\!\{g_{\leftrightarrow}\}$, where $g_{\leftrightarrow}$ denotes left-right reflection, defined by the axis of left-right symmetry $e_{\leftrightarrow}$. In this case, $H\!=\!\text{O(3)}/\{g_{\leftrightarrow}\}\!=\!\text{SO(3)}$.
@@ 4.3 Implementation details
We augment training volumes with commonly used spatial (rigid transformations, scaling, flips) and intensity (bias field, gamma, Gaussian noise, low resolution) transforms [perezgarcia2021, billot2023] in volumetric brain MRI. To enable our method to generalize to navigator volumes interleaved with diagnostic slices, we additionally simulate a spin history artifact from the preceding diagnostic slice [miller2015, gagoski2016], which appears as dark shading along the slice imaging plane (blue arrows in Fig. 3). We model this artifact with a Gaussian approximation of the slice profile:
@@ 5.1 Data
We evaluate the methods on Clinical-Young, a dataset of routine clinical volumes in 60 younger fetuses (GA 18-23 weeks). This dataset presents two challenges: clinical volumes exhibit a substantial domain shift from high-quality research data, and underdeveloped anatomy complicates pose estimation [zhan2013]. The voxel size is 1.76-3.5mm, and volumes are not interleaved with diagnostic slices. We also test the models on Navigators, a dataset containing 47 time-series (1210 total volumes) of 3D MRI navigator volumes interleaved with 2D diagnostic slices in 9 subjects (GA 26-36 weeks). Due to fast and interleaved acquisitions, these volumes have voxel size 4-6mm and contain real spin history artifacts. This dataset represents the largest domain gap from the research data, and volumes are representative of the intended clinical application, i.e, adaptive slice prescription. Both clinical datasets are acquired at the Boston Children’s Hospital.
@@ 5.4 Results
Figure 5: Simulation results. Left: Quantitative comparison of diagnostic slice stacks obtained using motion-blind prescription and E(3)-Pose. Mean $\pm$ standard deviation statistics are displayed. * indicates statistical significance (p<0.05p\!<\!0.05, pairwise Wilcoxon). Right: Brain coverage of the diagnostic slice stacks prescribed by each method, for three different example subjects and target anatomical orientations. Coverage gap (red) and obliqueness (purple, ) metrics are respectively displayed. Spatial coverage gap regions are outlined in red.
@@ Appendix A Automated slice prescription in fetal brain MRI
Figure 7: Interleaved acquisition framework. To correct for fetal head motion, we aim to automatically adjust the imaging plane $P_{k}$ of every diagnostic 2D slice, using the current head pose $T_{k}$. We rapidly acquire a navigator volume $f_{k}$ before every slice in order to accurately estimate $T_{k}$ from $f_{k}$ with E(3)-Pose.
@@
Figure 8: Real-time feedback loop system. Our implementation runs on a 3T Siemens scanner, connected to a server hosted on a GPU-enabled laptop. On the server (right), E(3)-Pose separately estimates the rotation $R_{k}$ and translation $t_{k}$ of the fetal head pose $T_{k}$, based on the navigator volume $f_{k}$ sent by the scanner (left). We apply translation $t_{k}$ to center the next navigator volume FOV around the fetal head. Future work will apply the full rigid transform $T_{k}$ to prescribe the next 2D diagnostic slice to account for the current head pose. Finally, we send the prescription parameters to the scanner for the next acquisition.
@@
We developed a real-time feedback loop system using existing software [wighton2024] that runs this sequence on a 3T Siemens scanner (Fig. 8). Specifically, we have implemented the translation estimation step of E(3)-Pose and used it to translate the FOV of every navigator volume on pregnant patients in real time. The videos in the linked project page demonstrate how the navigator volumes are automatically shifted to follow the translational movements of the fetal head. Future work will implement real-time prescription of 2D diagnostic slices, using the full pose $T_{k}$ returned by E(3)-Pose in every navigator volume.

### v4

@@ Abstract
Project page: ramyamut.github.io/e3-pose/ This work is supported by the the MIT Abdul Latif Jameel Clinic, the MIT-Takeda Program, the Chou Family Transformative Research Fund, NIH NIBIB (R01EB032708, 5T32EB1680), and NIH NICHD (R01HD100009). This work involved human subjects or animals in its research. Approval of all ethical and experimental procedures and protocols was granted by the Institutional Review Board of Boston Children’s Hospital.
@@ 1 Introduction
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [billard2019], autonomous driving [chiu2021], and image-guided surgery [yaniv2006]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology [glenn2009, gholipour2014] (Appendix A). Due to inter-slice fetal motion, oblique slices (i.e., misaligned with the target anatomical orientation) and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack (Fig. 1) [malamateniou2013, xu2023, white2010]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The pose estimated from each navigator volume can be used to adaptively adjust the imaging plane of the next 2D slice in real time (Fig. 1) [hess2011, tisdall2012, gagoski2016]. Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and low SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) 2D slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates a left-right symmetry, introducing pose ambiguities [zhan2013]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@ 4.2 Rotation estimation
Let $G_{\text{symm}}\!=\!\{g_{s}\in\text{O(3)}\!:\!f\!\circ\!g_{s}^{-1}\!=\!f\}$ be the object symmetry group [saxena2007]. To handle object symmetries, we aim to parametrize $h(R)$ such that $\rho_{h}(g_{s})h(R)\!=\!h(R)$ for all $g_{s}\!\in\!G_{\text{symm}}$ and $R\!\in\!\text{SO(3)}$; and $h$ remains continuous over the quotient group $H\!=\!\text{O(3)}/G_{\text{symm}}$. Here, we consider the case $G_{\text{symm}}\!=\!\{g_{\leftrightarrow}\}$, where $g_{\leftrightarrow}$ denotes left-right reflection, defined by the axis of left-right symmetry $e_{\leftrightarrow}$. In this case, $H\!=\!\text{O(3)}/\{g_{\leftrightarrow}\}\!=\!\text{SO(3)}$.
@@ 4.3 Implementation details
We augment training volumes with commonly used spatial (rigid transformations, scaling, flips) and intensity (bias field, gamma, Gaussian noise, low resolution) transforms [perezgarcia2021, billot2023] in volumetric brain MRI. To enable our method to generalize to navigator volumes interleaved with diagnostic slices, we additionally simulate a spin history artifact from the preceding slice [miller2015, gagoski2016], which appears as dark shading along the slice imaging plane (blue arrows in Fig. 3). We model this artifact with a Gaussian approximation of the slice profile:
@@ 5.1 Data
We evaluate the methods on Clinical-Young, a dataset of routine clinical volumes in 60 younger fetuses (GA 18-23 weeks). This dataset presents two challenges: clinical volumes exhibit a substantial domain shift from high-quality research data, and underdeveloped anatomy complicates pose estimation [zhan2013]. The voxel size is 1.76-3.5mm, and volumes are not interleaved with diagnostic slices. We also test the models on Navigators, a dataset containing 47 time-series (1210 total volumes) of 3D MRI navigator volumes interleaved with 2D slices in 9 subjects (GA 26-36 weeks). Due to fast and interleaved acquisitions, these volumes have voxel size 4-6mm and contain real spin history artifacts. This dataset represents the largest domain gap from the research data, and volumes are representative of the intended clinical application, i.e, adaptive slice prescription. Both clinical datasets are acquired at the Boston Children’s Hospital.
@@ 5.4 Results
Figure 5: Simulation results. Left: Quantitative comparison of diagnostic slice stacks obtained using motion-blind prescription and E(3)-Pose. Mean $\pm$ standard deviation statistics are displayed. * indicates statistical significance (p<0.05p\!<\!0.05, pairwise Wilcoxon). Right: Brain coverage of the slice stacks prescribed by each method, for three different example subjects and target anatomical orientations. Coverage gap (red) and obliqueness (purple, ) metrics are respectively displayed. Spatial coverage gap regions are outlined in red.
@@ Appendix A Automated slice prescription in fetal brain MRI
Figure 7: Interleaved acquisition framework. To correct for fetal head motion, we aim to automatically adjust the imaging plane $P_{k}$ of every 2D slice, using the current head pose $T_{k}$. We rapidly acquire a navigator volume $f_{k}$ before every slice in order to accurately estimate $T_{k}$ from $f_{k}$ with E(3)-Pose.
@@
Figure 8: Real-time feedback loop system. Our implementation runs on a 3T Siemens scanner, connected to a server hosted on a GPU-enabled laptop. On the server (right), E(3)-Pose separately estimates the rotation $R_{k}$ and translation $t_{k}$ of the fetal head pose $T_{k}$, based on the navigator volume $f_{k}$ sent by the scanner (left). We apply translation $t_{k}$ to center the next navigator volume FOV around the fetal head. Future work will apply the full rigid transform $T_{k}$ to prescribe the next slice to account for the current head pose. Finally, we send the prescription parameters to the scanner for the next acquisition.
@@
We developed a real-time feedback loop system using existing software [wighton2024] that runs this sequence on a 3T Siemens scanner (Fig. 8). Specifically, we have implemented the translation estimation step of E(3)-Pose and used it to translate the FOV of every navigator volume on pregnant patients in real time. The videos in the linked project page demonstrate how the navigator volumes are automatically shifted to follow the translational movements of the fetal head. Future work will implement real-time prescription of slices, using the full pose $T_{k}$ returned by E(3)-Pose in every navigator volume.